OpenStudy (anonymous):

Solve for x: 2|x − 3| + 1 = 7 A.) x = 0, x = 6 B.) x = −1, x = 6 C.) x = 0, x = 9 D.) No solutions

OpenStudy (anonymous):

Subtract 1 from both sides, then divide both sides by 2, you get /x-3/=3 so x- 3 = 3 and x- 3 = -3, therefore, x = 6 and x = 0

OpenStudy (anonymous):

There are two cases for which equation can be written and solved Case1: when x-3>0 .........|x-3| = x-3 and Case2: when x-3<0.......|x-3|= -(x-3)